OpenStudy (anonymous):

Solve the system of equations by elimination: x - 6y = -3 and 3x + 6y = 15 A. (3, -1) B. (3, 1) C. (1, 3) D. (1, -3)

jimthompson5910 (jim_thompson5910):

Add the equations to get x - 6y = -3 3x + 6y = 15 -------------- 4x + 0y = 12 4x = 12 x = ??

OpenStudy (anonymous):

are you still need help?

OpenStudy (anonymous):

Yes if it isn't c?

jimthompson5910 (jim_thompson5910):

if 4x = 12, then what is x?

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

4 times _____ = 12 what goes in the blank?

OpenStudy (anonymous):

4x4 = 12 ??

OpenStudy (anonymous):

That isn't there.

jimthompson5910 (jim_thompson5910):

don't worry about the answer choices right now 4 times SOME NUMBER= 12 what replaces "SOME NUMBER" ?

OpenStudy (anonymous):

1,3?

jimthompson5910 (jim_thompson5910):

I can see why you picked that, but it's incorrect

jimthompson5910 (jim_thompson5910):

x = 3 is the solution to 4x =12

jimthompson5910 (jim_thompson5910):

if x = 3, then what is y?

OpenStudy (anonymous):

B?

jimthompson5910 (jim_thompson5910):

good, the answer is (3,1)

OpenStudy (anonymous):

Thanks!